The cost of a machine is $400000.The machine can be used for 6 years.The salvage value of the machine is 25% of the cost of the machine.Tabulate the net depreciation and book value of the machine for each year by sinking fund method.The interest rate is 12% compounded annually
1
Expert's answer
2020-11-30T16:43:53-0500

Annual depreciation is "R = (400,000 - 100,000)*\frac{0.12}{1.12^6 - 1} = 36,967.72."

Book value at the end of 1st year = 400,000 - 36,967.72 = 363,032.28.